OpenStudy (anonymous):

Use the slope formula to find the missing coordinate. T(13, 2), W( X , -5) m =

OpenStudy (anonymous):

m=-7/2

OpenStudy (anonymous):

But what is the slope??

OpenStudy (anonymous):

\[\frac{13+5}{2-x}=-\frac{7}{2}\] \[\frac{18}{2-x}=-\frac{7}{2}\]or to make it easier try \[\frac{18}{x-2}=\frac{7}{2}\] solve the ratio for \(x\)

OpenStudy (anonymous):

\[36=7(x-2)\] etc

OpenStudy (anonymous):

I got -7 as my answer...

OpenStudy (anonymous):

Are you doing wrong??

OpenStudy (anonymous):

yeah my answer isn't working but neither is his,

OpenStudy (anonymous):

the answer is 15.

OpenStudy (anonymous):

I think now also you are wrong...

OpenStudy (anonymous):

it works in the calculator.

OpenStudy (anonymous):

\[36 = 7(x - 2)\] Now distribute 7 to right hand side, What you got tell me...

OpenStudy (anonymous):

15 was correct. :)